# Shared Spatial Memory Through Predictive Coding

### v1

@@ Abstract
Sharing and reconstructing a consistent spatial memory is a critical challenge in multi-agent systems, where partial observability and limited bandwidth often lead to catastrophic failures in coordination. We introduce a multi-agent predictive coding framework that formulate coordination as the minimization of mutual uncertainty among agents. Instantiated as an information bottleneck objective, it prompts agents to learn not only who and what to communicate but also when. At the foundation of this framework lies a grid-cell-like metric as internal spatial coding for self-localization, emerging spontaneously from self-supervised motion prediction. Building upon this internal spatial code, agents gradually develop a bandwidth-efficient communication mechanism and specialized neural populations that encode partners’ locations—an artificial analogue of hippocampal social place cells (SPCs). These social representations are further enacted by a hierarchical reinforcement learning policy that actively explores to reduce joint uncertainty. On the Memory-Maze benchmark, our approach shows exceptional resilience to bandwidth constraints: success degrades gracefully from 73.5% to 64.4% as bandwidth shrinks from 128 to 4 bits/step, whereas a full-broadcast baseline collapses from 67.6% to 28.6%. Our findings establish a theoretically principled and biologically plausible basis for how complex social representations emerge from a unified predictive drive, leading to social collective intelligence.
@@ Introduction
Collective intelligence is a cornerstone of biological success, enabling groups of organisms to perform complex tasks that far exceed the capabilities of any single individual [29, 41, 1, 7, 4]. From ant colonies forging optimal foraging trails to wolf packs executing intricate hunting strategies, the ability to form and act upon a shared understanding of the world is paramount [3, 36, 14]. This capacity is particularly salient in spatial navigation, where sociable animals such as bats in a cave complex or rats in a maze coordinate their movements by aligning internal representations of their environment [38, 33, 17]. Such alignment is often achieved through a sparse exchange of high-level cues, such as sonar chirps, ultrasonic squeaks, or visual signals, suggesting the existence of a shared cognitive map: a dynamic, distributed neural representation of the environment that includes the positions of resources, hazards, and, crucially, conspecifics [26, 45, 5, 46, 13].
@@
The neural algorithms underlying these shared representations are thought to be a specialized extension of the mechanisms supporting individual cognition. In mammals, the hippocampal-entorhinal system provides the substrate for an individual’s cognitive map, with place cells encoding specific locations and grid cells, providing a metric scaffold for space [18, 37, 16, 15]. The discovery of “social place cells”—neurons that fire when a partner is at a particular location—offers compelling evidence for a dedicated neural substrate that integrates the self with peers into this spatial framework [47, 40, 24]. These findings mean that the brain possesses sophisticated machinery not only for building a model of its own world but also for representing the world of others, a prerequisite for any meaningful social coordination. However, there is a critical distinction between representing others within a single brain and coordinating across brains. Within the centralized architecture of the hippocampal-entorhinal system, different neural populations can access and integrate social information through dense internal connectivity—communication is not a limiting factor. The computational challenge emerges at a different scale: when multiple individuals must align their internal spatial models through external communication that is severely bandwidth-constrained. This raises a fundamental question: What computational principles enable agents to develop socially-aware representations from individual experience and to coordinate these representations across a distributed system despite limited communication?
@@
Similarly, in artificial multi-agent systems, replicating this biological prowess faces a same persistent challenge: the “communication bottleneck” [8, 19, 30, 10]. While multi-agent systems hold immense promise for applications in exploration, search, and rescue, their collective performance often degrades catastrophically as communication bandwidth becomes a bottleneck [20, 34, 35]. From an information theoretic perspective, the core issue is the inefficient management of statistical redundancy. When several agents explore an environment concurrently, their sensory observations are highly correlated. Naive communication strategies, such as broadcasting raw or compressed sensory data, inevitably lead to channel bandwidth starvation when transmitting vast amounts of redundant information that a partner could likely infer on its own [9, 12, 43, 11]. This leaves a critical gap in our understanding: How can a decentralized system learn an efficient communication mechanism that is not only sparse, but also optimally informative, when transmitting only the essential information needed to resolve a partner’s uncertainty?
@@
In this work, we demonstrate that shared spatial memory can emerge from a single computational principle: minimizing mutual predictive uncertainty between agents. We develop a multi-agent predictive coding framework that implements this principle through three integrated levels: Level 1 (Individual Perception)—spontaneous formation of grid-cell-like spatial metrics for self-localization; Level 2 (Social Communication)—bandwidth-efficient communication mechanisms and emergent social place cell representations; and Level 3 (Strategic Exploration)—hierarchical policies that reduce collective uncertainty through coordinated exploration.
To instantiate this principle, we ground our approach in the synthesis of predictive coding and information theory, constructing a robust shared spatial memory system that operates under extreme communication constraints. Our approach, summarized in Fig. 1a-c, addresses a cooperative navigation scenario where multiple agents leverage egocentric vision to explore an environment and interact via emergent symbols (Fig. 1a). The framework is designed to first enable each agent to individually build a predictive, allocentric bird’s-eye-view (BEV) map from its own observations (Fig. 1b). Critically, it then provides a mechanism for agents to exchange these predictive insights through an information bottleneck, a process that minimizes their partners’ future uncertainty and construct an efficient, socially grounded shared representation (Fig. 1c). Our framework is realized through two tightly integrated mechanisms. First, each agent individually learns to perceive its environment by building a predictive model that generates a BEV map from egocentric vision input (Fig. 1b). We show that this complex visuospatial inference task is critically scaffolded by an internal path integrator that spontaneously self-organizes to produce a stable, grid-cell-like representation, providing a consistent metric for space without any explicit supervision (Fig. 2). This finding reinforces the idea that a robust internal model of self-motion is a foundational prerequisite for coherent world-building. Second, building upon this perceptual foundation, agents collaboratively develop a communication mechanism (Fig. 1c). Rather than exchanging raw observations, agents learn to transmit compressed, discrete symbols that are optimally tailored to reduce their partners’ future uncertainty. This communication mechanism emerges from a variational information bottleneck (VIB) objective, which provides a principled trade-off between minimizing communication cost (rate) and maximizing predictive utility for a partner (distortion) [22, 12]. The fundamental observation is that agents can learn how to efficiently communicate with each other due to more interactions and hence sharing more prior information, which is similar to the fundamental principle of emergent communications[32, 31]. The communication mechanism is embedded within a hierarchical reinforcement learning framework (termed HRL-ICM), where an ICM-aided MAPPO algorithm makes strategic navigation decisions to efficiently reduce uncertainty [42, 28].
@@
The proposed HRL-ICM framework achieves a level of performance and robustness that substantially exceeds current baselines such as No Communication, Periodic Broadcast, and Full Broadcast. As validated in the challenging Memory-Maze benchmark [27], our framework exhibits remarkable resilience to communication constraints. While the success rate of the Full Broadcast baseline collapses from 67.6% to 28.6% (a 58% relative decline) when the bandwidth is reduced from 128 to 4 bits/step, our method degrades slightly from 73.5% to 64.4% (only a 12% relative decline), maintaining superior performance even under extreme bandwidth constraints and underscoring the efficiency of the learned predictive communication mechanism (Fig. 5g). Most remarkably, we discover that the objective of predicting a partner’s state drives the emergence of a functionally specialized neural substrate for social cognition. Within the network’s social decoding module, neurons spontaneously segregate into distinct populations, including units that selectively encode the location of specific teammates—a striking artificial analogue of social place cells observed in the mammalian hippocampus (Fig. 4) [21]. Causal analyses confirm that these emergent representations are not epiphenomenal but functionally critical for effective coordination (Fig. 6b and Supplementary Video 3).
@@
By unifying principles from predictive coding and information bottleneck theory [38, 6], our work provides a theoretically grounded and biologically explainable basis for shared spatial memory. The main contributions of this work are threefold. ❶ We develop a novel multi-agent framework that enables an efficient, semantically rich communication mechanism to emerge from first principles. ❷ We provide a plausible computational model for the function and emergence of social place cells, suggesting they arise as a necessary component of a social predictive coding system. ❸ We validate this framework in a challenging benchmark, demonstrating state-of-the-art performance, scalability, and bandwidth insensitivity. This work paves the way for a new generation of collaborative agent systems that can coordinate with the efficiency and flexibility of their biological counterparts, grounded in a deeper understanding of the computational principles that may underlie collective intelligence itself.
@@
Figure 2: Grid-cell-like representations enhance robust global BEV mapping. a, 2D spatial firing rate maps of learned grid-like representations. b, Spatial autocorrelograms of representative grid-like neurons reveal hexagonal symmetry. c, The convergence of loss functions during gird cell network training. d, As training proceeds, top neuron gridness increases while path integration error decreases. e, Ablation study shows the full model achieves lower trajectory error and higher prediction confidence than the variant without the grid scaffold. f, Comparison of BEV map reconstructed by different configurations.
@@ Results
A fundamental prerequisite for our predictive coding framework is an agent’s ability to form a stable internal model of its own state and surroundings. As outlined in Fig. 1b, this is achieved by solving two coupled prediction problems: predicting the visual appearance of the world (BEV mapping) and predicting the agent’s own trajectory through it (path integration). This section presents results for this first component of our unified framework—the individual predictive model—demonstrating that solving the self-supervised task of continuous self-motion prediction naturally reinstates a spatial coding scheme, the reminiscent of biological grid cells, which in turn provides an essential metric scaffold for robust visual perception.
@@
The core of the path integration module is an LSTM network, tasked with predicting its future pose based only on its past velocity commands. Under this predictive constraint, the network’s hidden units spontaneously form periodic spatial firing patterns similar to the grid-cell representations previously reported in both biological and artificial agents [2, 37]. Many individual units developed highly structured spatial firing fields arranged in a triangular lattice (Fig. 2a). To quantitatively confirm this observation, we computed the spatial autocorrelogram for each unit’s activation map. This analysis revealed a clear hexagonal symmetry in the firing patterns of key neurons (Fig. 2b), the defining physiological signature of grid cells. The mathematical formulation for calculating these rate maps, autocorrelograms, and the final gridness score is detailed in Supplementary Method S1.1.3.
@@
This grid-like representation is not only a byproduct but also a computational regime that the network converges to when optimizing for stable self-motion prediction. To demonstrate this, we track the co-evolution of this neural code and the module’s predictive performance throughout training. As the network learned, the gridness score of the most prominent units—a measure of hexagonal symmetry—steadily increased, while the path integration error, which quantifies the error in the predicted trajectory, concurrently and dramatically decreased (Fig. 2d and Supplementary Fig. S4). This tight correlation, along with the steady convergence of the training losses (Fig. 2c), demonstrates that the formation of a stable, periodic neural code is the very mechanism through which the network masters the task of motion prediction. The network learns to represent space metrically, enabling it to accurately integrate its movements over long distances and maintain a coherent belief about its location.
@@
The primary function of this grid-like scaffold is to stabilize the agent’s visual prediction process. By conditioning the BEV generation model on the latent state provided by the path integrator, the agent can correctly register and align transient visual inputs into a coherent allocentric frame. To isolate and quantify its functional role, we performed an ablation study, comparing the full predictive model against a variant where the grid-cell scaffold was disabled. The scaffold’s contribution is clear: the full model achieved significantly lower trajectory error and higher prediction confidence (Fig. 2e). This improvement in self-localization directly translates into superior visual prediction. BEV maps from the scaffolded model are more complete and geometrically accurate, while the no-scaffold baseline products fragmented and distorted maps (Fig. 2f). Supplementary Fig. S6 visualizes this full pipeline, showing how egocentric views are converted to BEV predictions and integrated into a persistent spatial memory. Supplementary Fig. S7 illustrates the progressive and cumulative construction of the shared spatial memory, visualizing how the map is built up at sequential time intervals. Further examples of neural activations consistently show a transformation from disorganized firing to highly regular, grid-like fields (Supplementary Fig. S4). The dynamic evolution of this process is visualized in Supplementary Video 5, which shows the transformation from random initialization to highly structured grid-cell-like representations across training epochs. Therefore, these findings show that the emergence of grid-cell-like coding in our predictive model serves a necessary functional role for constructing a stable spatial memory, providing a computational foundation for coherent world modeling.
@@
Figure 3: An efficient, structured, and intelligent communication mechanism emerges from a predictive objective. a. Intelligent communication strategies emerge, with message frequency (heatmaps) concentrated at critical decision points like coordination hubs or dead ends, demonstrating strategic triggering. b. The emergent protocol is highly bandwidth-efficient, consistently requiring the lowest communication overhead across diverse maze types when compared to full and periodic broadcast baselines. c. The protocol is theoretically controllable via the information bottleneck’s $\beta$ coefficient, which enables a principled trade-off between message compression (compression ratio) and predictive utility (reconstruction accuracy). d. An emergent symbolic vocabulary is grounded in strategic contexts. A t-SNE visualization reveals distinct symbol clusters corresponding to high-level navigational situations, such as encountering a “Three-way Fork” or discovering the “Target”. e. Communication causally influences decision-making. In a controlled scenario where one agent faces a choice between an unexplored path (A) and a known one (B), communication from its partner allows it to identify Path A as the more informative route. f. The behavioral impact is statistically significant. A violin plot quantifying choices at two-way forks shows that communication leads to a significantly higher probability of selecting the unexplored, informative path compared to the no-communication baseline (***p < 0.001), confirming the protocol’s causal role in improving efficiency.
Having established that individual agents can build robust predictive models of their environment, we next investigate how the framework extends this principle to the social domain. Agents learn to cooperate by developing a communication mechanism guided by a singular objective: transmit only the information that maximally reduces a partner’s future uncertainty. This pressure compels the agents to collaboratively discover a communication system that is not only sparse but also intelligent and semantically structured.
@@
A key property of the emergent protocol is its context-aware transmission strategy. Agents learn to communicate not periodically, but strategically at particular moments and locations where a partner’s internal model is most likely to be inaccurate. We visualize the spatial distribution of communication events and find that agents concentrate their transmissions at points of high predictive uncertainty (Fig. 3a). For example, in mazes with a central hub, communication peaks in this area, where an agent’s next move is most ambiguous. In such locations, information from others is crucial to avoid redundant exploration. Conversely, in mazes with long, deceptive dead ends, agents learn to communicate most frequently from deep within these traps. This behavior is a direct solution to the social prediction problem: a message from a dead end serves as a powerful “prediction error” signal to teammates, effectively correcting their erroneous implicit prediction that the path might be fruitful. This strategic triggering, consistently observed across diverse maze topologies (Supplementary Fig. S7), demonstrates that agents learn an implicit model of their partners’ beliefs, sharing information precisely when it can best resolve uncertainty and prevent predictive mistakes.
This strategy of transmitting only the most surprising, uncertainty-reducing information naturally gives rise to a highly bandwidth-efficient protocol. By eliminating the transmission of predictable, redundant information, the system minimizes its communication overhead. We quantify this by comparing the average number of messages per episode against standard baselines. Across thousands of randomly generated layouts, our framework consistently operates with a fraction of the bandwidth required by periodic or full-broadcast approaches (Fig. 3b), confirming that the social predictive objective is a powerful first principle for learning efficient communication.
Furthermore, the protocol is fundamentally structured and controllable, adhering to the principles of information bottleneck theory that underpin our social predictive model. The information bottleneck’s $\beta$ coefficient allows for a principled tuning of the trade-off between compression (rate) and predictive utility (distortion). As we increase $\beta$, placing a stronger penalty on information rate, the agents are forced to develop a more compressed, abstract symbolic representation (Fig. 3c). This comes at the expected cost of reduced predictive utility, confirming that the framework provides theoretical control over the communication channel’s properties, allowing it to be adaptable to varying bandwidth constraints.
@@
The uncertainty-driven compression scheme culminates in a meaningful symbolic vocabulary. A t-SNE visualization of the message latent space reveals that symbols are not used randomly but form distinct clusters corresponding to high-level strategic contexts (Fig. 3d). For instance, specific symbol families emerge for situations such as navigating a “Three-way Fork”, encountering a “Dead-end”, or discovering the “Target”. These are all contexts that demand clear communication to resolve ambiguity. To test for a causal link between these symbols and agent behavior, we conduct a controlled experiment (Fig. 3e). An agent at a fork could resolve its uncertainty about which path to choose only after receiving a message from its partner. We quantify these choices across thousands of trials and find that communication significantly and reliably guides agents to select the more informative, unexplored path over a known one (Fig. 3f). This provides compelling evidence that the social predictive objective forges a functional and intelligent communication system. The agents do not merely signal raw data; they learn to transmit compressed, symbolic representations of prediction errors to collaboratively refine their shared world model.
@@
Figure 4: Predictive learning forges a functionally specialized social place code. a, Model architecture. Observer ($S_{1}$) and partner ($S_{2}$) states are processed through a bottleneck layer and relational head. The network is trained by back-propagating predictive error from self, partner, and social outputs. b, Functionally distinct neuron types. Heatmaps show spatial firing fields for representative “Pure SPCs” (tuned to partner location), “Pure Place Cells” (tuned to self location), and “Special SPCs” (mixed selectivity). c, Population code for inter-agent distance. Top panels show 2D maps and 1D tuning curves for neurons selective for close-, mid-, and far-distances. Bottom right, a heatmap of all distance-tuned neurons reveals a “tiling” of the distance space. d, Quantitative functional dissociation. Top, bar plots show high mutual information (MI) with self-position for Place Cells and with partner-position for SPCs. Bottom, a scatter plot of self MI vs. partner MI reveals specialized cell clusters. e, Causal necessity of SPCs demonstrated via in-silico lesioning. Targeted lesioning of distance-selective SPCs (“SPC Lesion”, left) specifically impairs distance prediction, while lesioning the entire peer LSTM module (right) impairs general position prediction, compared to controls. f, Co-evolution of performance and specialization. Left, trajectory predictions are accurate for the trained network but poor for untrained or SPC-lesioned networks. Right, validation loss decreases over training epochs as the proportion of specialized Place Cells and SPCs increases.
A central challenge in collective intelligence is the capability to form and maintain representations of others. An agent must move beyond a purely egocentric worldview and model the states, locations, and potential actions of its partners. We hypothesize that the predictive objective of forecasting a teammate’s future sensory state would provide a sufficiently rich learning signal to induce a functionally specialized neural substrate for social spatial cognition. To test this, we design a social processing module that integrates the agent’s own state ($S_{1}$) with information about its partner’s state ($S_{2}$), and uses the resulting joint representation to predict future outcomes for both agents (Fig. 4a). Analysis of the learned internal representations revealed clear functional specialization with properties consistent with vector- and grid-like spatial coding in artificial agents [47, 40, 24].
@@
Single-unit analyses indicate a spontaneous segregation into distinct, interpretable cell types (Fig. 4b). One substantial population behaved like classical place cells (network units encoding self-position, exhibiting sharp and stable firing fields tuned exclusively to the agent’s own location. These units are largely invariant to the partner’s motion or position, thereby providing a stable allocentric representation of self. In contrast, a second major population fired as a function of the partner’s location: these artificial social place cells (SPCs; units encoding partner position, distinct from the self-position-encoding place cells) showed strong spatial tuning to the partner’s location within the observer’s reference frame (Fig. 4b, Neurons 29 and 36), providing a substrate for tracking others. We also observe a population of mixed-selectivity units that conjunctively encode self- and partner-locations (Fig. 4b, Neuron 8). Representative galleries across four conditions (self-moving, partner-moving, both-moving, both-static) indicate that this specialization is expressed across units and task contingencies (Supplementary Figs. S8-S10 and Video 2).

### v2

@@ Abstract
Constructing a consistent shared spatial memory is a critical challenge in multi-agent systems, where partial observability and limited bandwidth often lead to catastrophic failures in coordination. We introduce a multi-agent predictive coding framework that formulates coordination as the minimization of mutual uncertainty among agents. Through an information bottleneck objective, this framework prompts agents to learn not only who and what to communicate but also when. At the foundation of this framework lies a grid-cell-like metric as internal spatial coding for self-localization, emerging spontaneously from self-supervised motion prediction. Building upon this internal spatial code, agents gradually develop a bandwidth-efficient communication mechanism and specialized neural populations that encode partners’ locations—an artificial analogue of hippocampal social place cells (SPCs). These social representations are further utilized by a hierarchical reinforcement learning policy that actively explores to reduce joint uncertainty. On the Memory-Maze benchmark, our approach shows exceptional resilience to bandwidth constraints: success degrades gracefully from 73.5% to 64.4% as bandwidth shrinks from 128 to 4 bits/step, whereas a full-broadcast baseline collapses from 67.6% to 28.6%. Our findings establish a theoretically principled and biologically plausible basis for how complex social representations emerge from a unified predictive drive, leading to collective intelligence.
@@ Introduction
Collective intelligence enables groups to perform complex tasks exceeding individual capabilities [1, 2, 3, 4, 5, 6]. From ant colonies to wolf packs, forming and acting upon shared environmental understanding is paramount [7, 8, 9]. In spatial navigation, animals such as bats or rats coordinate movements by aligning internal spatial representations [10, 11, 12]. This alignment emerges through sparse exchange of high-level cues (sonar chirps, ultrasonic squeaks, visual signals), suggesting the existence of shared cognitive maps: distributed neural representations of environments including resources, hazards, and conspecifics [13, 14, 15, 16, 17].
@@
Neural algorithms underlying shared representations extend innate mechanisms supporting individual cognition. In mammals, the hippocampal-entorhinal system provides the substrate for cognitive maps, with place cells encoding locations and grid cells providing a metric scaffold [18, 19, 20, 21, 22]. “Social place cells”—neurons firing when a partner occupies particular locations—offer compelling evidence for neural substrates integrating self with peers [23, 24, 25]. These findings indicate that the brain possesses sophisticated machinery not only for building a model of its own world but also for representing the world of others, a prerequisite for any meaningful social coordination. However, a critical distinction exists between representing others within single brains versus coordinating across brains through severely bandwidth-constrained communication. This raises a fundamental question: What computational principles enable agents to develop socially-aware representations and to coordinate these representations across distributed systems under limited communication?
@@
Artificial multi-agent systems face the same challenge: the “communication bottleneck” [26, 27, 28, 29]. While promising for exploration, search, and rescue, collective performance degrades catastrophically when bandwidth is limited [30, 31, 32]. The core issue is inefficient management of statistical redundancy among agents. Concurrent exploration yields highly correlated sensory observations; naive strategies (broadcasting raw or compressed data) waste bandwidth transmitting redundant information partners could infer [33, 34, 35, 36]. How can decentralized systems learn a sparse yet optimally informative communication mechanism, transmitting only essential information to resolve partner uncertainty?
@@
We demonstrate that shared spatial memory emerges from minimizing mutual predictive uncertainty between agents. Our multi-agent predictive coding framework implements this through: Level 1—grid-cell-like spatial metrics for self-localization; Level 2—bandwidth-efficient communication and emergent social place cells; Level 3—hierarchical policies reducing collective uncertainty through coordinated exploration.
We ground our approach in predictive coding and information theory, constructing a shared spatial memory system operating under strict communication constraints. Agents with egocentric vision explore environments and communicate via emergent symbols (Fig. 1a). Each agent builds predictive bird’s-eye-view (BEV) maps from observations (Fig. 1b), then exchanges insights through an information bottleneck minimizing partners’ uncertainty and constructing efficient shared representations. Visualized in Fig. 1c, networks develop neurons tuned to self-location, partner location, or both, mirroring the mammalian hippocampus (Fig. 4b). The framework integrates two mechanisms. First, agents build predictive models generating BEV maps from vision, scaffolded by internal path integrators spontaneously producing grid-cell-like representations without supervision (Fig. 2). Second, agents develop communication mechanisms (Fig. 3), transmitting compressed symbols that reduce partners’ uncertainty. This emerges from variational information bottleneck (VIB) objectives balancing communication cost and predictive utility [37, 34, 38, 39], embedded in hierarchical reinforcement learning with intrinsic curiosity (HRL-ICM) for strategic uncertainty reduction [40, 41].
@@
Our HRL-ICM framework substantially exceeds baselines (No Communication, Periodic Broadcast, Full Broadcast). On Memory-Maze [42], it exhibits remarkable resilience: while Full Broadcast collapses from 67.6% to 28.6% (58% relative decline) when bandwidth is reduced from 128 to 4 bits/step, our method degrades minimally from 73.5% to 64.4% (12% decline), maintaining superior performance under extreme constraints (Fig. 5g). Predicting partners’ states drives emergence of specialized social cognition substrates. Neurons spontaneously segregate into populations selectively encoding teammates’ locations—artificial social place cells mirroring mammalian hippocampus (Fig. 4) [43]. Causal analyses confirm these representations are functionally critical (Fig. 6b, Supplementary Video 3).
@@
Unifying predictive coding and information bottleneck theory [10, 44], our work provides a theoretically grounded, biologically plausible basis for shared spatial memory. Our main contributions are as follows: ❶ A multi-agent framework enabling efficient, semantically rich communication from first principles. ❷ A computational model for social place cell emergence as necessary components of social predictive coding. ❸ Validation demonstrating state-of-the-art performance, scalability, and bandwidth insensitivity. This work enables collaborative systems coordinating with biological efficiency, grounded in computational principles underlying collective intelligence.
@@
Figure 2: Grid-cell-like representations enhance robust global BEV mapping. a, 2D spatial firing rate maps of learned grid-like representations. Each panel shows activity of a single unit in the path integrator LSTM network, labeled by artificial neuron index (#) and gridness score ($G_{60}$, range: -2 to +2; higher values indicate stronger hexagonal symmetry). Heat maps represent normalized firing rates across the 2D environment. b, Spatial autocorrelograms (SACs) of representative grid-like units reveal hexagonal symmetry. Each circular plot shows the autocorrelation structure of the corresponding unit’s firing pattern from panel a, with the six-fold rotational symmetry characteristic of biological grid cells. Unit labels and $G_{60}$ scores match panel a. c, The convergence of loss functions during grid cell network training. d, As training proceeds, top unit gridness ($G_{60}$) increases while path integration error decreases, demonstrating the co-emergence of structured representations and predictive accuracy. e, Ablation study shows the full model achieves lower trajectory error and higher prediction confidence than the variant without the grid scaffold. f, Comparison of BEV map reconstructed by different configurations.
@@ Results
A fundamental prerequisite for our predictive coding framework is an agent’s ability to form a stable internal model of its own state and surroundings. As outlined in Fig. 1b, this is achieved by solving two coupled prediction problems: predicting the visual appearance of the world (BEV mapping) and predicting the agent’s own trajectory through it (path integration). While previous work has demonstrated grid-cell-like representations can emerge in navigation tasks [45, 19], these studies primarily focused on supervised learning with explicit spatial labels. Here, we demonstrate that self-supervised self-motion prediction alone—without any explicit spatial supervision—naturally develops a grid-cell-like spatial coding scheme that provides an essential metric scaffold for robust visual perception.
@@
The path integration module uses an LSTM network to predict its future pose based solely on past velocity commands. Under this predictive constraint, the network’s hidden units spontaneously form periodic spatial firing patterns. Many individual units developed highly structured spatial firing fields arranged in triangular lattices (Fig. 2a), labeled by neuron index and gridness score ($G_{60}$)—a quantitative measure of hexagonal symmetry ranging from -2 to +2. To verify this grid-like organization quantitatively, we computed spatial autocorrelograms (SACs) for each neuron’s activation map. This analysis revealed clear hexagonal symmetry in firing patterns (Fig. 2b), the defining physiological signature of grid cells found in mammalian entorhinal cortex [18].
@@
Importantly, this grid-like representation is not merely a byproduct but represents a computational solution that the network actively converges to when optimizing for stable self-motion prediction. To demonstrate this, we tracked the co-evolution of this neural code and predictive performance throughout training. As the network learned, gridness scores of the most prominent units steadily increased, while path integration error concurrently and dramatically decreased (Fig. 2c,d and Supplementary Fig. S4). This tight correlation demonstrates that the formation of stable, periodic neural coding is the very mechanism through which the network masters motion prediction, enabling it to accurately integrate movements over long distances and maintain coherent beliefs about its location.
@@
Beyond characterizing the emergence of grid-like codes, we establish their functional necessity for visual prediction. By conditioning BEV generation on the latent state provided by the path integrator, agents can correctly register and align transient visual inputs into coherent allocentric frames. To isolate and quantify this functional role, we performed ablation studies comparing the full predictive model against variants where the grid-cell scaffold was disabled. The scaffold’s contribution is clear: the full model achieved significantly lower trajectory error and higher prediction confidence (Fig. 2e). This improvement in self-localization directly translates into superior visual prediction. BEV maps from scaffolded models are more complete and geometrically accurate, while no-scaffold baselines produce fragmented and distorted maps (Fig. 2f). Thus, these findings demonstrate that the emergence of grid-cell-like coding in our predictive model serves a necessary functional role for constructing stable spatial memory, providing a computational foundation for coherent world modeling.
@@
[…]
Having established that individual agents can build robust predictive models of their environment through grid-cell-like scaffolds, we next investigate how the framework extends this predictive principle to the social domain. A central challenge in multi-agent coordination is learning efficient communication under bandwidth constraints. While existing approaches rely on handcrafted communication protocols [26, 27] or learn to broadcast raw sensory data [30, 31], these methods struggle with bandwidth limitations and fail to discover semantically meaningful communication strategies. Unlike frameworks exploring asymmetric, teacher–student knowledge transfer [46], our setting addresses symmetric, peer-to-peer coordination. We compare our HRL-ICM framework against three baselines: No Communication, where agents search independently; Periodic Broadcast, where agents communicate at fixed intervals under bandwidth constraints; and Full Broadcast, where agents exchange information without bandwidth limitations. Framing communication through the information bottleneck principle, we demonstrate that agents can learn to cooperate by developing a communication mechanism guided by a singular objective: transmit only information that maximally reduces a partner’s future uncertainty. This simple requirement compels agents to collaboratively discover an emergent communication mechanism that is not only sparse and bandwidth-efficient, but also intelligent and semantically structured.
@@
A key property of the emergent mechanism is its context-aware, task-oriented transmission strategy. Unlike prior work where agents communicate periodically or in response to predefined triggers, our agents learn to communicate strategically at particular moments and locations where a partner’s internal model is most likely to be inaccurate. We visualize the spatial distribution of communication events and find that agents concentrate their transmissions at points of high predictive uncertainty (Fig. 3a). For example, in mazes with a central hub, communication peaks in this critical coordination area where an agent’s next move is most ambiguous. Conversely, in mazes with long, deceptive dead ends, agents learn to communicate most frequently from deep within these traps. This behavior directly solves the social prediction problem: a message from a dead end serves as a powerful “prediction error” signal to teammates, effectively correcting their erroneous implicit prediction that the path might be fruitful. This strategic triggering, consistently observed across diverse maze topologies (Supplementary Fig. S7), demonstrates that agents learn an implicit model of their partners’ beliefs, sharing information precisely when it can best resolve uncertainty and prevent predictive mistakes.
Building upon this context-aware triggering mechanism, the strategy of transmitting only the most surprising, uncertainty-reducing information naturally gives rise to a highly bandwidth-efficient communication mechanism. Across thousands of randomly generated layouts, our framework consistently operates with a fraction of the bandwidth required by periodic or full-broadcast approaches (Fig. 3b). Furthermore, the mechanism is fundamentally structured and controllable through information bottleneck theory. The $\beta$ coefficient enables principled tuning between source compression and predictive utility. Increasing $\beta$ forces more compressed, abstract symbolic representations (Fig. 3c), confirming that our framework provides theoretical control over communication effectiveness.
@@
Most importantly, the uncertainty-driven compression scheme culminates in a meaningful symbolic vocabulary. A t-SNE visualization of the message latent space reveals distinct clusters corresponding to high-level strategic contexts: “Three-way Fork”, “Dead-end”, “Target” (Fig. 3d). To test for a causal link between these symbols and agent behavior, we conduct controlled experiments where an agent at a fork could resolve its uncertainty only after receiving a message from its partner (Fig. 3e). Quantifying choices across thousands of trials, we find that communication significantly guides agents to select the more informative, unexplored path over a known one (Fig. 3f), providing compelling evidence that agents learn to transmit compressed, symbolic representations of prediction errors to collaboratively refine their shared world model and reduce collective uncertainty.
@@
Figure 4: Predictive learning forges a functionally specialized social place code. a, Model architecture. Observer ($S_{1}$) and partner ($S_{2}$) states are processed through a bottleneck layer and relational head. The network is trained by back-propagating predictive error from self, partner, and social outputs. b, Functionally distinct neuron types. Three artificial neuron types in the relational head: Pure Place Cells (self-location encoding, Unit 40), Pure SPCs (partner-location encoding, Units 29, 36), and Special SPCs (mixed selectivity, Unit 8). Heatmaps show normalized firing rates. c, Population code for inter-agent distance. Top panels show 2D maps and 1D tuning curves for neurons selective for close-, mid-, and far-distances. Bottom right, a heatmap of all distance-tuned neurons reveals a “tiling” of the distance space. d, Quantitative functional dissociation. Top, bar plots show high mutual information (MI) with self-position for Place Cells and with partner-position for SPCs. Bottom, a scatter plot of self MI vs. partner MI reveals specialized cell clusters. e, Causal necessity of SPCs demonstrated via in-silico lesioning. Left, targeted lesioning of relational-layer distance-selective SPCs ($n=20$ units) specifically impairs distance prediction compared to intact model and random lesion controls (two-tailed unpaired $t$-test; $n=50$ test episodes per condition). Right, targeted lesioning of LSTM-layer SPCs impairs position prediction. Violin plots show full data distribution; *$p<0.05$, **$p<0.01$, ***$p<0.001$. f, Co-evolution of performance and specialization. Left, trajectory predictions are accurate for the trained network but poor for untrained or SPC-lesioned networks. Right, validation loss decreases over training epochs as the proportion of specialized Place Cells and SPCs increases.
In the previous section, we demonstrated that agents learn efficient communication through predictive coding. However, a more fundamental question remains: What internal representations must agents develop to effectively model and predict the states of their partners? While prior work has identified place cells encoding self-location in artificial agents [45, cueva2018emergence], the neural substrates for representing social partners in spatial contexts remain largely unexplored. Here, we show that a central challenge in collective intelligence—forming and maintaining representations of others—can be addressed through predictive learning.
@@
We design a social processing module integrating self-state ($S_{1}$) with partner-state ($S_{2}$) to predict future outcomes for both agents (Fig. 4a). Single-unit analyses indicate spontaneous segregation into distinct, interpretable cell types (Fig. 4b). One substantial population behaved like classical place cells, exhibiting sharp and stable firing fields tuned exclusively to the agent’s own location. These units are largely invariant to the partner’s motion or position, thereby providing a stable allocentric representation of self. In contrast, a second major population fired as a function of the partner’s location: these artificial social place cells (SPCs) showed strong spatial tuning to the partner’s location within the observer’s reference frame (Fig. 4b, Neurons 29 and 36), providing a substrate for tracking others. We also observe a population of mixed-selectivity units that conjunctively encode self- and partner-locations (Fig. 4b, Neuron 8). Representative galleries across four conditions (self-moving, partner-moving, both-moving, both-static) indicate that this specialization is expressed across units and task contingencies (Supplementary Figs. S8-S10).